Simplify and express each of the following in exponential form:
 (2^3 × 2)^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

= (2^3+1 )^2 ... [∵a^m × a^n = a^m+n ]
= (2^4)^2
= (2^4)^2 can be written as = (2)4x2 ... [∵(a^m)^n= a^mn ]
= 2^8

Thanks